# Discovering Message Passing Hierarchies for Mesh-Based Physics Simulation

### v1

@@ 5 Related Work
Learning-based physics simulation.
Recent literature has shown that learning-based simulators can efficiently handle complex and high-dimensional problems, such as fluid dynamics (Zhu et al., 2024), structural analysis (Kavvas et al., 2018; Thai, 2022), and climate modeling (Kurth et al., 2018; Rasp et al., 2018; Rolnick et al., 2022). The models can be roughly categorized into three groups based on data representation: those modeling partial differential equations (Raissi et al., 2017; 2019; Lu et al., 2019; Li et al., 2021; Wang et al., 2021), particle-based systems (Li et al., 2019; Sanchez-Gonzalez et al., 2020; Ummenhofer et al., 2020; Prantl et al., 2022), and mesh-based systems (Pfaff et al., 2021; Lino et al., 2022; Fortunato et al., 2022; Cao et al., 2023). The rapid inference time and differentiable property of these models greatly facilitate downstream tasks, such as inverse design (Wang & Zhang, 2021; Goodrich et al., 2021; Allen et al., 2022; Janny et al., 2023).
GNN-based physics simulation.
Previous work has explored GNNs in various physical domains, such as articulated systems (Sanchez-Gonzalez et al., 2018), soft-body deformation and fluids (Li et al., 2019; Mrowca et al., 2018; Sanchez-Gonzalez et al., 2020; Rubanova et al., 2022; Wu et al., 2023), rigid body dynamics (Battaglia et al., 2016; Li et al., 2019; Mrowca et al., 2018; Bear et al., 2021; Rubanova et al., 2022), and aerodynamics (Belbute-Peres et al., 2020; Hines & Bekemeyer, 2023; Pfaff et al., 2021; Fortunato et al., 2022; Cao et al., 2023). Among them, MGN (Pfaff et al., 2021) is a key method that models mesh-based dynamics through graph interactions. Subsequent approaches primarily focus on enhancing modeling capabilities and reducing computational costs.
Hierarchical GNNs for physics simulation.
Hierarchical GNNs employ multi-scale graph structures (Lino et al., 2022; Han et al., 2022; Fortunato et al., 2022; Allen et al., 2023; Janny et al., 2023; Cao et al., 2023; Yu et al., 2024) to decrease overhead by using fewer nodes at coarser levels and enabling long-range feature propagation. GMR-Transformer-GMUS (Han et al., 2022) employs a uniform sampling pooling method to select pivotal nodes. MS-MGN (Fortunato et al., 2022) uses a dual-level hierarchical GNN and performs message passing at both fine and coarse resolutions. Hierarchical GNNs with multi-level structures Lino et al. (2022); Cao et al. (2023); Yu et al. (2024) are most relevant to our approach, as they integrate message passing neural networks within the U-Net architecture (Ronneberger et al., 2015). Lino et al.(Lino et al., 2022) uses manually set grid resolutions and spatial proximity for graph pooling, which requires predefined parameters. BSMS-GNN (Cao et al., 2023) introduces a bi-stride pooling strategy that pools nodes on alternating breadth-first search frontiers while enhancing edges with two-hop connections. HCMT (Yu et al., 2024) refines the structure further by applying Delaunay triangulation to bi-stride nodes. Notably, these methods construct multi-level structures as preprocessing and cannot change the graph hierarchies under varying physical conditions. Moreover, they typically use uniform feature aggregation for intra-level propagation, which may hinder the directed transfer of significant dynamic patterns, or use attention-based aggregation, which increases computational overhead. Furthermore, inter-level propagation is often predefined or unlearnable, limiting flexibility in transferring information across hierarchy levels. In contrast, our model generates context-aware and temporally evolving graph hierarchies and incorporates learnable anisotropic feature propagation, allowing it to better adapt to various initial conditions and rapidly changing dynamic systems.
@@ 6 Conclusions and Limitations
In this paper, we introduced DHMP, a neural network that significantly advances the state-of-the-art in mesh-based physics simulation. Our key innovation is dynamically creating the context-aware graph structures of hierarchical GNNs through a differentiable node selection process. To this end, we proposed an anisotropic message passing mechanism to enhance the propagation of long-term dependencies between distant nodes, aligning with the directed nature of significant dynamic patterns. Extensive experiments show that DHMP outperforms existing models, especially those with fixed graph hierarchies, in both short-term and long-term predictions.
A potential limitation of this work is the need to improve the interpretability of the learned hierarchy structure. Additionally, we would consider incorporating specific physical priors into DHMP to further enhance the model’s robustness and generalizability, particularly in resolution-free problem settings, which have been less explored in existing mesh-based approaches.
Ethics Statement
In this work, we adhere to the highest ethical standards across all stages of research. No human subjects were involved, and no personal data was used, ensuring compliance with privacy and security protocols. All datasets utilized are publicly available, mitigating concerns related to sensitive information exposure. We acknowledge the potential use of physics simulation models for harmful insights if misapplied; therefore, we encourage careful consideration of the context and application domain when deploying these models.
Reproducibility Statement
We mainly build DHMP based on the released code of BSMS-GNN. We prioritize the repeatability of our work and will open source the source code. All results can be reproduced by following the experimental details presented in Section 4.1 and Appendices AC.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62250062, 62106144), the Shanghai Municipal Science and Technology Major Project (Grant No. 2021SHZDZX0102), and the Fundamental Research Funds for the Central Universities.
Appendix A Dataset
We utilize four established datasets from MGN (Pfaff et al., 2021): CylinderFlow, Airfoil, Flag, and DeformingPlate.
  • The CylinderFlow case examines the transient incompressible flow field around a fixed cylinder positioned at different locations.
@@ Appendix A Dataset
The CylinderFlow, Airfoil, and Flag datasets are each split into $1000$ training sequences, $100$ validation sequences, and $100$ testing sequences. The DeformingPlate dataset is split into $500$ training sequences, $100$ validation sequences, and $100$ testing sequences.
We also consider a more challenging dataset FoldingPaper where of varying forces at the four corners drags paper with time-varying lagrangian mesh graphs, generated using the ARCSim solver (Narain et al., 2012; Wu et al., 2023). This dataset is split into $500$ training sequences, $100$ validation sequences, and $100$ testing sequences.
The statistics of five datasets are shown in Table 6. We also list the input physical quantities in Table 7.
Table 6: Statistics of the CylinderFlow, Airfoil, Flag, DeformingPlate and FoldingPaper datasets.
Dataset Average # nodes Average # edges Mesh type # Hierachies # Steps
@@
Table 7: Comparisions of the edge offsets and node inputs of different physical systems.
@@ Appendix B Model Implementation
We list model configurations of different physical systems below.
@@
  • Edge offsets. $X$ and $x$ stand for the mesh-space and world-space position. For an Eulerian system, only mesh position is used for $\mathbf{e}_{ij}$, while for a Lagrangian system, both mesh-space and world space position are used. The edge offsets are directly used as low-dimensional input to the edge update function $\phi^{e}$. This means that these features are concatenated together and fed directly into $\phi^{e}$ without first passing through an MLP or other encoding process to transform them into a higher-dimensional representation.
@@
  • Input and target of the physical term of node $v_{i}$. $v$ is the velocity, $\rho$ is the density, $P$ is the absolute pressure, and the dot $\dot{a}=a_{t+1}-a_{t}$ stands for temporal change for a variable $a$. $\dot{n}$ stands for the node type of $v_{i}$. Random Gaussian noise is added to the node input features to enhance robustness during training (Pfaff et al., 2021; Sanchez-Gonzalez et al., 2020; Cao et al., 2023). All the variables involved are normalized to zero-mean and unit variance via preprocessing. The preprocessed physical term is fed to the encoder to transform it into a high-dimensional representation.
@@
The encoder, decoder, node update function $\phi^{v}$, and edge update function $\phi^{e}$ all utilize two-layer MLPs with ReLU activation and a hidden size of $128$. Similarly, the importance weight network $\phi^{w}$ in AMP is implemented using a two-layer MLP. LayerNorm is applied to the MLP outputs, except for the decoder and the importance weight network. We set $K=2$ for edge enhancement, which is aligned with the setting of BSMS-GNN (Cao et al., 2023). In the Gumbel-Softmax for differentiable node selection, temperature annealing decreases the temperature from $5$ to $0.1$ using a decay factor of $\gamma=0.999$, encouraging exploration of hierarchies while gradually refining their selection to ensure stability. DHMP is trained with Adam optimizer, using an exponential learning rate decay from $10^{-4}$ to $10^{-6}$. All experiments are conducted using 4 Nvidia RTX 3090. We mainly build DHMP based on the released code of BSMS-GNN (Cao et al., 2023).
Appendix C Baseline Details
We compare DHMP with four competitive baselines: (1) MGN (Pfaff et al., 2021) which performs multiple message passing on the input high-resolution mesh topology; (2) BSMS-GNN (Cao et al., 2023), which uses predefined bi-stride pooling prior as preprocessing to generate static hierarchies on same mesh topology and ; (3) Lino et al.(Lino et al., 2022), which uses manually set grid resolutions and spatial proximity for graph pooling; (4) HCMT (Yu et al., 2024), which uses Delauny triangulation based on bi-stride nodes and adopt attention mechanism to enable non-uniform feature propagation. All models are trained with Adam optimizer, using an exponential learning rate decay from $10^{-4}$ to $10^{-6}$.
  • MGN  We apply 15 iterations of message passing in all datasets. The encoder, decoder, node update function and edge update function are set up the same way as in our model.
  • BSMS-GNN  We use the same levels of hierarchy as DHMP for BSMS-GNN for preprocessing data. We use the minimum average distance as the seeding heuristic for BFS search recommended in its original paper. The multi-level building is processed in one pass. The inter-level propagation uses the normalized node degree to convolve features from neighbors to central nodes. The encoder, decoder, node update function and edge update function are set up the same way as in our model.
  • Lino et al.  We use the four-scale GNN structure proposed in Lino et al. (2022). The edge length of the smallest cell for each dataset is 1/10 of the average scene size, with each lower scale doubling in size. Each of the message-passing networks has two hidden layers per MLP.
  • HCMT  We set the number of blocks in HCMT to $15$. The hidden dimension and the heads are set as $128$ and $4$. We use the same hierarchy levels as DHMP on all datasets. For the Cylinder and Airfoil datasets, since there are hollow sections in the mesh, we did not use Delaunay triangulation for remeshing. Instead, we utilized edge connections generated by bi-stride pooling.
    Appendix D Additional Results
    D.1 Ablation Study
@@
    In Sec. 4.3, we compare different variants of our DHMP model against the BSMS-GNN baseline, to evaluate the effectiveness of (i) dynamic hierarchy construction based on the input mesh topology and physical quantities, (ii) anisotropic intra-level feature propagation, (iii) learnable inter-level feature propagation. The variants we investigate include:
    • Static-Anisotropic-Unlearnable (M1): (ii),
    • Static-Anisotropic-Learnable (M2): (ii+ iii),
    • Dynamic-Anisotropic-Unlearnable (M3): (i)+(ii).
@@
    In this ablation study, we utilize a static graph hierarchy preprocessed using bi-stride pooling as described in the BSMS-GNN paper (Cao et al., 2023), along with a non-parametric intra-level aggregation function from previous works (Pfaff et al., 2021; Cao et al., 2023). Additionally, BSMS-GNN employs unlearnable node degree metrics to generate inter-level aggregation weights, which convolve features based on the normalized node degree for inter-level propagation. We show the quantitative RMSE values of Figure 4 in Table 8.
    Table 8: Qualitative results for model variants of DHMP and baselines on one-step and long-term prediction errors.
@@
    Static-Anisotropic-Unlearnable (M1) 0.1995 0.4804 9.621 121.1
    Static-Anisotropic-Learnable (M2) 0.1695 0.4666 8.317 109.9
@@
    DHMP 0.1568 0.3049 6.571 76.16
    D.2 Edge Enhancement
    When constructing the lower-level graph $\mathcal{G}_{l+1}$ based on the selected nodes, the edges $\mathcal{E}_{l+1}$ are formed by connecting these nodes using the original edges $\mathcal{E}_{l}$ from the previous graph. However, this approach may lead to disconnected partitions, as observed in previous studies (Lee et al., 2019; Cao et al., 2023; Gao & Ji, 2019), and illustrated in Figure 7. To address this issue, we enhance the connectivity of $\mathcal{E}_{l+1}$ by incorporating $K$-hop edges during the edge construction process. We investigate the impact of different $K$ values, specifically $K=2,3,4$, on the Flag dataset. The results are presented in Table 9.

### v2

@@ 5 Related Work
Learning-based and GNN-based physical simulation.
Recent advances have demonstrated that learning-based approaches can efficiently tackle complex and high-dimensional physical simulation tasks, including fluid dynamics (Zhu et al., 2024), structural analysis (Kavvas et al., 2018; Thai, 2022), and climate modeling (Kurth et al., 2018; Rasp et al., 2018; Rolnick et al., 2022; Lam et al., 2023). These methods can be broadly categorized by their data representations: partial differential equations (Raissi et al., 2017, 2019; Lu et al., 2019; Li et al., 2021; Wang et al., 2021), particle-based systems (Li et al., 2019; Sanchez-Gonzalez et al., 2020; Ummenhofer et al., 2020; Prantl et al., 2022), and mesh-based systems (Pfaff et al., 2021; Lino et al., 2022; Fortunato et al., 2022; Cao et al., 2023). The rapid inference and differentiable nature of these models have facilitated a range of downstream applications, such as inverse design (Wang & Zhang, 2021; Goodrich et al., 2021; Allen et al., 2022; Janny et al., 2023). In particular, Graph Neural Networks (GNNs) have emerged as a powerful tool for modeling physical systems across various domains, including articulated bodies (Sanchez-Gonzalez et al., 2018), soft-body deformation and fluids (Li et al., 2019; Mrowca et al., 2018; Sanchez-Gonzalez et al., 2020; Rubanova et al., 2022; Wu et al., 2023), rigid body dynamics (Battaglia et al., 2016; Li et al., 2019; Mrowca et al., 2018; Bear et al., 2021; Rubanova et al., 2022), and aerodynamics (Belbute-Peres et al., 2020; Hines & Bekemeyer, 2023; Pfaff et al., 2021; Fortunato et al., 2022; Cao et al., 2023). Among these, MeshGraphNets (Pfaff et al., 2021) serves as a representative, introducing a general scheme for representing meshes as graphs and learning mesh-based dynamics, inspiring subsequent works that focus on improving modeling capacity and computational efficiency.
Hierarchical GNNs for physical simulation.
Hierarchical GNNs leverage multi-scale graph structures (Lino et al., 2022; Han et al., 2022; Fortunato et al., 2022; Allen et al., 2023; Janny et al., 2023; Cao et al., 2023; Yu et al., 2024) to reduce computational overhead by operating on coarser representations and to facilitate long-range information propagation. For example, GMR-Transformer-GMUS (Han et al., 2022) employs uniform sampling for pooling, while Eagle (Janny et al., 2023) adopts a two-scale message passing scheme with geometric clustering by preprocessing the mesh structure. LayersNet (Shao et al., 2023) introduces a static, patch-based hierarchy for garment animation, simplifying interaction modeling via particle patches. More recent works (Lino et al., 2022; Cao et al., 2023; Yu et al., 2024; Garnier et al., 2024; Hy & Kondor, 2023) integrate hierarchical GNNs with U-Net architectures (Ronneberger et al., 2015), using static multi-level structures for message passing. For instance, Lino et al. (2022) relies on manually defined grid resolutions, BSMS-GNN (Cao et al., 2023) proposes a bi-stride pooling strategy with enhanced edge connectivity, and HCMT (Yu et al., 2024) further refines the hierarchy using Delaunay triangulation and replaces message passing with graph attention. However, these methods typically employ precomputed, static hierarchies and uniform feature aggregation, limiting their adaptability to dynamic physical environments. In contrast, our approach constructs context-aware, temporally evolving graph hierarchies with learnable anisotropic feature propagation, enabling robust adaptation to diverse initial conditions and rapidly changing dynamics.
Differentiable graph pooling.
A variety of differentiable and learnable graph pooling methods have been proposed to enable end-to-end training of hierarchical graph representations, such as DiffPool (Ying et al., 2018), TopKPool (Gao & Ji, 2019), SAGPool (Lee et al., 2019), and ASAPooling (Ranjan et al., 2020). These methods construct hierarchical representations of graphs by either learning soft cluster assignments or selecting top-ranked nodes based on learned importance scores. However, most of these methods are primarily designed for static graphs and focus on global graph-level tasks, where unpooling or reconstruction of the original graph structure is not required. In contrast, mesh-based physical simulation demands fine-grained local information and temporally-evolving hierarchies for accurate modeling of physical dynamics. Our approach differs by constructing context-aware, time-varying graph hierarchies tailored for mesh-based simulation, enabling flexible integration of global and local features and supporting dynamic adaptation to changing physical conditions.
@@ 6 Conclusions and Limitations
In this paper, we introduced EvoMesh, a neural network that significantly advances the state-of-the-art in mesh-based simulation. Our key innovation is adaptively creating the temporally-evolving and context-aware graph structures of hierarchical GNNs through a differentiable node selection process. To this end, we proposed an anisotropic message passing mechanism to enhance the propagation of long-term dependencies between distant nodes, aligning with the directed nature of significant dynamic patterns. Extensive experiments show that EvoMesh outperforms existing models, especially those with fixed graph hierarchies, in both short-term and long-term predictions.
A potential limitation of this work is the need to improve the interpretability of the learned hierarchies. Additionally, we would consider incorporating physical priors into EvoMesh to enhance the model’s robustness and generalizability, particularly in resolution-free problem settings, which have been less explored in existing mesh-based approaches.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (62250062), the Smart Grid National Science and Technology Major Project (2024ZD0801200), the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), and the Fundamental Research Funds for the Central Universities.
Impact Statement
In this work, we adhere to the highest ethical standards throughout all stages of research. No human subjects were involved, and no personal data was used, ensuring full compliance with privacy and security protocols. All datasets employed are publicly available, minimizing concerns regarding sensitive information exposure.
We recognize that while physical simulation models can drive significant advancements, they also have the potential for misuse if applied irresponsibly. Thus, we emphasize the importance of thoughtful consideration regarding the context and ethical implications when deploying these models, particularly in high-stakes applications such as healthcare, engineering, and environmental management.
Appendix
Appendix A Datasets
We employ four established datasets from MGN (Pfaff et al., 2021): CylinderFlow, Airfoil, Flag, and DeformingPlate.
  • The CylinderFlow case examines the transient incompressible flow field around a fixed cylinder positioned at different locations, with varying inflow velocities.
@@ Appendix A Datasets
The CylinderFlow, Airfoil, and Flag datasets are each split into $1{,}000$ training sequences, $100$ validation sequences, and $100$ testing sequences. The DeformingPlate dataset is split into $500$ training sequences, $100$ validation sequences, and $100$ testing sequences.
We also consider a more challenging dataset, FoldingPaper, where varying forces at the four corners deform paper with time-varying Lagrangian mesh graphs, generated using the ARCSim solver (Narain et al., 2012; Wu et al., 2023). This dataset is divided into $500$ training sequences, $100$ validation sequences, and $100$ testing sequences.
We present statistical details of all five datasets in Table 6 and the input physical quantities in Table 7.
Table 6: Statistics of the CylinderFlow, Airfoil, Flag, DeformingPlate, and FoldingPaper datasets.
Dataset Average # nodes Average # edges Mesh type # Hierarchies # Steps
@@
Table 7: Comparisons of the edge offsets and node inputs of different physical systems.
@@ Appendix B Model Implementation
We present model configurations of different physical systems below:
@@
  • Edge offsets. $X$ and $x$ stand for the mesh-space and world-space position. For an Eulerian system, only mesh position is used for $\mathbf{e}_{ij}$, while for a Lagrangian system, both mesh-space and world-space positions are used. The edge offsets are directly used as low-dimensional input to the edge update function $\phi^{e}$. These features are concatenated and fed into $\phi^{e}$ without any transformation through an MLP or other encoding processes to generate a higher-dimensional representation.
@@
  • Input and target of the physical term of node $v_{i}$. $v$ is the velocity, $\rho$ is the density, $P$ is the absolute pressure, and the dot $\dot{a}=a_{t+1}-a_{t}$ stands for temporal change for a variable $a$. ${n}$ stands for the node type of $v_{i}$. Random Gaussian noise is added to the node input features to enhance robustness during training (Pfaff et al., 2021; Sanchez-Gonzalez et al., 2020; Cao et al., 2023). All the variables involved are normalized to zero-mean and unit variance via preprocessing. The preprocessed physical term is fed to the encoder to transform it into a high-dimensional representation.
@@
The encoder, decoder, node update function $\phi^{v}$, and edge update function $\phi^{e}$ all utilize two-layer MLPs with ReLU activation and a hidden size of $128$. Similarly, the importance weight network $\phi^{w}$ in AMP is implemented using a two-layer MLP. LayerNorm is applied to the MLP outputs, except for the decoder and the importance weight network. We set $K=2$ for edge enhancement, which is aligned with the setting of BSMS-GNN (Cao et al., 2023). In the Gumbel-Softmax for differentiable node selection, temperature annealing decreases the temperature from $5$ to $0.1$ using a decay factor of $\gamma=0.999$, encouraging exploration of hierarchies while gradually refining their selection to ensure stability. EvoMesh is trained with Adam optimizer, using an exponential learning rate decay from $10^{-4}$ to $10^{-6}$. All experiments are conducted using 4 Nvidia RTX 3090. We mainly build EvoMesh based on the released code of BSMS-GNN (Cao et al., 2023).
Appendix C Additional Results
C.1 Ablation Study
@@
In Sec. 4.3, we compare different variants of our EvoMesh model against the BSMS-GNN baseline, to evaluate the effectiveness of (i) dynamic hierarchy construction based on the input mesh topology and physical quantities, (ii) anisotropic intra-level feature propagation, (iii) learnable inter-level feature propagation. The variants we investigate include:
  • Static(Bi-stride)-Anisotropic-Unlearnable (M1),
  • Static(Bi-stride)-Anisotropic-Learnable (M2),
  • Uniform-Anisotropic-Learnable (M3),
  • Dynamic-Anisotropic-Unlearnable (M4).
@@
In this ablation study, we utilize a static graph hierarchy preprocessed using bi-stride pooling as described in the BSMS-GNN paper (Cao et al., 2023) for static hierarchy variants, along with a non-parametric intra-level aggregation function from previous works (Pfaff et al., 2021; Cao et al., 2023). Additionally, BSMS-GNN employs unlearnable node degree metrics to generate inter-level aggregation weights, which convolve features based on the normalized node degree for inter-level propagation. We show the quantitative RMSE values of Figure 4 in Table 8.
Table 8: Qualitative results of model variants of EvoMesh and the baseline model.
@@
Static(Bi-stride)-Anisotropic-Unlearnable (M1) 0.1995 0.4804 9.621 121.1
Static(Bi-stride)-Anisotropic-Learnable (M2) 0.1695 0.4666 8.317 109.9
Uniform-Anisotropic-Learnable (M3) 0.2019 0.3999 9.357 145.27
@@
EvoMesh 0.1568 0.3049 6.571 76.16
C.2 Alternative Implementation of AMP
We evaluate the effectiveness of the AMP module by comparing it with alternative attention-based designs (Veličković et al., 2018; Vaswani et al., 2017).
Unlike standard attention mechanisms such as Graph Attention Networks(GAT) (Veličković et al., 2018), AMP enables differentiable dynamic hierarchy construction—an essential capability for modeling evolving physical systems. While AMP and GAT both compute importance weights, their usage differs substantially: GAT uses these weights to aggregate node features, whereas AMP directly applies the predicted weights to edge features and reuses them for inter-level feature propagation. This dual mechanism allows AMP to jointly support dynamic hierarchy learning and physics modeling via end-to-end training. To assess the benefits of AMP over GAT, we replace AMP with GATConv (using a single attention head) in EvoMesh. As shown in Table 9, the AMP-based model consistently outperforms the GAT-based variant across datasets, highlighting the importance of dynamic hierarchy construction.
We further explore the flexibility of AMP by implementing it with a cross-attention mechanism (Vaswani et al., 2017) in place of the default MLP. Specifically, edge features $\hat{\mathbf{e}}_{ij}$ are refined by attending to node features $\mathbf{v}_{i}$ through scaled dot-product attention:
$\mathbf{\hat{e}}_{\text{aggr}}=\operatorname{CrossAttention}(Q=\mathbf{v}_{i},,K={\hat{\mathbf{e}}_{ij}},,V={\hat{\mathbf{e}}_{ij}}),$ (9)
and the resulting attention scores are reused for inter-level propagation. Experimental results in Table 9 show that while the cross-attention implementation is competitive, the original MLP-based AMP achieves better overall accuracy. This suggests that the edge features $\hat{\mathbf{e}}_{ij}$ already encode relevant node information, making the explicit inclusion of $\mathbf{v}_{i}$ through attention redundant and potentially less efficient.
Table 9: Results of different implementations of AMP module.
Cylinder Flag
AMP Implementation RMSE-1 (×$\times 10^{-2}$) RMSE-All (×$\times 10^{-2}$) RMSE-1(×$\times 10^{-2}$) RMSE-All (×$\times 10^{-2}$)
EvoMesh-GATConv 0.2025 10.253 0.3009 106.2
EvoMesh-CrossAttention 0.1881 8.356 0.3650 75.12
EvoMesh 0.1568 6.571 0.3049 76.16
C.3 Edge Enhancement
When constructing the lower-level graph $\mathcal{G}_{l+1}$ based on the selected nodes, the edges $\mathcal{E}_{l+1}$ are formed by connecting these nodes using the original edges $\mathcal{E}_{l}$ from the previous graph. However, this approach may lead to disconnected partitions, as observed in previous work (Lee et al., 2019; Cao et al., 2023; Gao & Ji, 2019), and illustrated in Figure 6. To address this issue, we enhance the connectivity of $\mathcal{E}_{l+1}$ by incorporating $K$-hop edges during the edge construction process. We investigate the impact of different $K$ values, specifically $K=2,3,4$, on the Flag dataset. The results are presented in Table 10,along with comparisons of the computational efficiency.